i can never figure out word problems can you please help. thank you in advance
108. Boxing match. A boxing ring is in the shape of a square, 20 feet on each side. How far apart are the fighters when they are in opposite corners of the ring?

Answer by checkley71(8403)

THE DIAGONAL OF A SQUARE 20 FOOT RING IS SOLVED BY THE FORMULA A^2+B^2=C^2 WHERE THE A & B SIDES ARE 20 FEET. THUS THE DIAGONAL (C) IS
20^2+20^2=C^2
400+400=C^2
800=C^2
C=SQRT800
C=28.28 FEET IS THE DIAGONAL ACROSS THE RING